(2/x) / (3x/4)= x/8
How do you solve this problem?

Answer by sujatha.krishna(11)

(2/x) / (3x/4)= x/8
2/x * 4/3x = x/8
8/3x^2 = x/8
Multiply the whole equation by 8
3x^2/8 = 8/x
3x^2 = 64/x
x^3 = 64/3
x = cube root (64/3)
x = 4/cube root(3)
x= 4/1.44
x = 2.77